OpenStudy (anonymous):

Help please? Which polynomial has a graph that passes through the given points? (–3, 10), (–1, 0), (0, –5), (3, 52)

OpenStudy (anonymous):

any clue as to the degree?

OpenStudy (anonymous):

unfortunatly no :/ thats all the information the question gives

OpenStudy (anonymous):

ok you have 4 points, so it can determine a polynomial of degree 3 since that has 4 coefficients

OpenStudy (anonymous):

since \(p(0)=-5\) youi know that the constant is \(-5\)

OpenStudy (anonymous):

y = x3 – 4x2 + 2x – 5 y = x3 + 4x2 – 2x – 5 y = x3 + 4x2 + 2x – 5 y = x4 + 4x3 – 2x2 – 5x here are the possible answers though

OpenStudy (anonymous):

ok since you have answers, you have a rather easier job check which numbers work

OpenStudy (anonymous):

last one is wrong because if you replace \(x\) by 0 you are supposed to get -5 you can cross that off the list

OpenStudy (anonymous):

@amistre64 has a nice way to do this

OpenStudy (amistre64):

xint yint (–3, 10), (–1, 0), (0, –5), (3, 52) high low lower high

OpenStudy (amistre64):

with multiple points its doable, but labor intensive

OpenStudy (amistre64):

it not the last one, since that y int is zero

OpenStudy (anonymous):

and you can replace \(x\) by \(-1\) and see which of the top three gives 0

OpenStudy (amistre64):

(–3, 10), (–1, 0), (0, –5), (3, 52) y = a + b(x+3) + c(x+3)(x+1) + d(x+3)(x+1) x

OpenStudy (anonymous):

actually i think that gives it right away

OpenStudy (amistre64):

yeah, it simpler to just fill in for x and do the math to narrow the options for this one

OpenStudy (anonymous):

i still would like to see your method what are \(a,b,c,d\)?

OpenStudy (anonymous):

looks like \(a=10\) from what you wrote

OpenStudy (anonymous):

which confuses me because the constant is \(-5\)

OpenStudy (anonymous):

oops i see \(a\) is not the constant i will shut up and learn

OpenStudy (amistre64):

lol, fine (–3, 10) y = a + b(x+3) + c(x+3)(x+1) + d(x+3)(x+1) x 10 = 10 0 0 0 y = 10 + b(x+3) + c(x+3)(x+1) + d(x+3)(x+1) x (–1, 0) y -10 = b(x+3) + c(x+3)(x+1) + d(x+3)(x+1) x 0 2b 0 0 -5 (0, –5) y = 10 -5(x+3) + c(x+3)(x+1) + d(x+3)(x+1) x y-10 = -5(x+3) + c(x+3)(x+1) + d(x+3)(x+1) x -5-10 -15 3c 0 0 (3, 52) y = 10 -5(x+3) + d(x+3)(x+1) x 52 = 10 - 30 d 6 4 3 72 = 72 d 1 y = 10 -5(x+3) + (x+3)(x+1) x

OpenStudy (anonymous):

oh!! replace \(x\) by -3 set result equal to 10 then move on down the line!

OpenStudy (amistre64):

correct, i learned later that this is a newton method, so that when more information is added, you dont have to redo all the stuff prior to it

OpenStudy (anonymous):

that is great!! also rather easier than solving a 4 by 4 system of equations. does this have a name?

OpenStudy (anonymous):

not "newton's method" as in "newton ralphson" right? must be something else

OpenStudy (amistre64):

legendre polynomial or some such is what i think we determined; i just call it the amistre64 chain of pain lol

OpenStudy (anonymous):

reminds me of pfd

OpenStudy (amistre64):

y = 10 -5(x+3) + (x+3)(x+1) x y = 10 - 5x - 15 + x^3 +x^2 +3x^2 +3x y = x^3 +4x^2 -2x -5

OpenStudy (anonymous):

damn i can't get a trademark symbol in latex any ideas?

OpenStudy (amistre64):

... dont use latex ;)

OpenStudy (anonymous):

i don't do windows

OpenStudy (amistre64):

^{tm} maybe?

OpenStudy (amistre64):

\[amistre64^{tm}\]

OpenStudy (amistre64):

\[\textsuperscript{\texttrademark}\] doesnt seem to be coded

OpenStudy (anonymous):

Thanks guys! Great explanations :D